OpenStudy (anonymous):

Tides are cyclical phenomena caused by the gravitational pull of the sun and the moon. On a particular retaining wall, the ocean generally reaches the 3 m mark at high tide. At low tide, the water reaches the 1 m mark. Assume that high tide occurs at 12:00 p.m. and at 12:00 a.m., and that low tide occurs at 6:00 p.m. and 6:00 a.m. What is the height of the water at 10:30 a.m.?

OpenStudy (anonymous):

is this correct 12pm to 12 am=12hrs 12pm to 10:30 am= 10.5 hrs h=(10.5/2)sm=2.625

OpenStudy (dumbcow):

what is "sm" ? i think the fact that they are cyclical implies they model a wave function like sine or cosine draw it out:

OpenStudy (dumbcow):

 looks like the cosine function with amplitude of 1 and period of 12 hours --> 2pi/b = 12, --> b = pi/6 y = cos(pi/6 x) + 2 plug in x = 10.5 y = 2.707